# Recent developments and research needs in turbulence modeling of hypersonic flows

### v1

@@ VII.2.1 Towards reliable data-driven models
Machine learning augmentations are generally developed as a functional map over a feature space. The underlying feature selection must balance both classical machine learning criteria and domain-specific requirements. From a classical machine learning perspective, features should provide unique and complementary information to prevent overfitting, even when dealing with correlated features. A parsimonious feature set is crucial, as it helps minimize overfitting risk, improves computational efficiency, and simplifies model interpretation. Features must also demonstrate robustness by being resistant to noise and small data variations, ensuring model stability and generalizability. The computational cost of calculating features must be considered, especially when working with large datasets.
@@
Generalization remains one of the most significant challenges in data-driven turbulence modeling. Model consistency is crucial for generalization, requiring careful consideration of the implicit relationship between state variables and augmentation values. Generalizable results in physical space can be achieved through careful interpolation over a well-designed feature space. However, this presents an important trade-off: while more features allow for more unique physical description, a lower-dimensional feature space is necessary for generalization with limited data. Model generalization to flows involving diverse physical processes is an active field of research, even for low-speed flows (Fang et al., 2023; Rincón et al., 2023; Cherroud et al., 2023; Oulghelou et al., 2024).

### v2

@@ VII.2.1 Towards reliable data-driven models
Machine learning augmentations are generally developed as a functional map over a feature space. The underlying feature selection must balance both classical machine learning criteria and domain-specific requirements. From a classical machine learning perspective, features should provide unique and complementary information to prevent overfitting, even when dealing with correlated features. A parsimonious feature set is crucial, as it helps minimize overfitting risk, improves computational efficiency, and simplifies model interpretation. Features must also demonstrate robustness by being resistant to noise and small data variations, ensuring model stability and generalizability. Feature sets have been proposed by various authors Ling and Templeton (2015); Wu, Xiao, and Paterson (2018), mostly in the context of low speed flows. More effort is required to develop principles and frameworks to systematically derive the features. For hypersonic flows, these are required not only to augment the Reynolds tensor, but also the unclosed terms in the energy equation, and in any additional transport equations,e.g., for reacting species. Of course, the computational cost of calculating the features must be considered, especially when working with large datasets. Approaches favoring sparsity, as well as feature selection strategies based on information theory Beneddine (2023) may help extracting optimal feature sets.
@@
Generalization remains one of the most significant challenges in data-driven turbulence modeling. Model consistency is crucial for generalization, requiring careful consideration of the implicit relationship between state variables and augmentation values. Generalizable results in physical space can be achieved through careful interpolation over a well-designed feature space. However, this presents an important trade-off: while more features allow for more unique physical description, a lower-dimensional feature space is necessary for generalization with limited data. In such cases, combining smaller models learned from subsets of data is empirically observed to be more efficient than training a single large model covering a wide range of flow processes Zhou (2025). Boosting, stacking, weighting, and voting approaches are examples of well-established ensemble approaches in modern machine learning. Model generalization to flows involving diverse physical processes is an active field of research, even for low-speed flows (Fang et al., 2023; Rincón et al., 2023; Cherroud et al., 2025; Oulghelou et al., 2024), and is warranted more efforts in the future.
Finally, machine learning methods often lack of interpretability. This is not only the case for black-box methods based on neural-network or random-forest approximations, but also for some symbolic models characterized by very complex expressions. Using sparsity-promoting techniques(Schmelzer, Dwight, and Cinnella, 2020) can help obtaining simpler models that are easier to interpret. Nonlinear sensitivity analysis tools, such as Shapley factors are another avenue for better understanding the learned corrections(He et al., 2022). More in general, explainable artificial intelligence (XAI) is a promising field with high potential impact in fluid flow studiesCremades et al. (2024).